OpenStudy (anonymous):

Let z be a complex number where z doesn't equal +/- i. Prove that z/(1+z^2) is real if and only if z is real or |z| = 1. I have an idea on how to solve this, but I want to know if the conjugate of z/(1+z^2) is (1+(conjugate of z)^2)/(conjugate of z)

OpenStudy (anonymous):

is that the proper conjugate?

OpenStudy (anonymous):

because I know that in order for it to be real z/(1+z^2) = (z/(1+z^2))*

OpenStudy (jamesj):

[ z/(1+z^2) ]* = z*/(1+z*^2)

OpenStudy (jamesj):

Actually your idea is a smart one. A complex number is real if and only if it equally its conjugate, yes.

OpenStudy (anonymous):

alright I will make an attempt with this

OpenStudy (jamesj):

Yes, this is a very nice approach.

OpenStudy (anonymous):

things are getting very large when I set z = (x+yi) can anything meaningful be obtained by keeping it as z and z* or should I keep going with this expansion?

OpenStudy (zarkon):

if z=0 then both sides are satisfied...so assume \(z\ne 0\) then if \[\frac{z}{1+z^2}\] is real then \[\frac{1+z^2}{z}=\frac{1}{z}+z\] is real

OpenStudy (zarkon):

let \[z=a+bi\] then you get \[\frac{a}{a^2+b^2}+a+\left(b-\frac{b}{a^2+b^2}\right)i\] this is real if b=0 or \(|z|=1\)

OpenStudy (anonymous):

what did you put the z =a + bi into to get that?

OpenStudy (zarkon):

\[\frac{1}{z}+z\]

OpenStudy (anonymous):

I'm finding how you got there hard to follow.

OpenStudy (anonymous):

wait I think I see where you multiplied the bottom by the conjugate

OpenStudy (anonymous):

at the beginning when you say both sides are satisfied what are you referring to? also does this satisfy an iff?

OpenStudy (anonymous):

alright I see what you are referring to now

OpenStudy (zarkon):

I'm saying if z=0 then z/(1+z^2) is real iff z is real

OpenStudy (zarkon):

I need to get that case out of the way since I reciprocated the function

OpenStudy (anonymous):

does this satisfy as the proof for an if and only if or do I need to show another case?

OpenStudy (zarkon):

what I wrote above is really an iff...so yes

OpenStudy (anonymous):

I find it amazing how you can just transform every problem into something so simple to interpret

OpenStudy (anonymous):

thanks a ton for the help on this one as well

OpenStudy (zarkon):

Not every problem ... I wish I could :)

OpenStudy (anonymous):

it just blows my mind every time you do it lol seems so complicated until you apply the zarkon magic

OpenStudy (jamesj):

Here's an alternative. If z/(z^2 + 1) is equal to its conjugate then \[ \frac{z^*}{1+z^{*2}} = \frac{z}{1+z^2} \] i.e., \[ z^* + z^2z^* = z + z^{*2}z \] i.e., \[ z^* - z = z^{*2}z - z^2z^* = ( z^* - z)zz^* \] Now that is true if and only \[ z^* - z = 0 \ \ \ or \ \ \ zz^* = 1 \] i.e., \[ z = z^* \ \ \ or \ \ \ |z|^2 = 1 \] i.e., either z is real or |z|^2 = 1

OpenStudy (zarkon):

nice solution

OpenStudy (jamesj):

Yes, it's a nice example of using the 'conjugate property' of real numbers.

OpenStudy (anonymous):

why does z* - z need to be 0?

OpenStudy (anonymous):

oh wait I see it

OpenStudy (jamesj):

z* - z = (z* - z)(zz*) hence zz* = 1, or if not, it must be that z* - z = 0. I.e., if a = ab, then b = 1 or a = 0

OpenStudy (anonymous):

that is a really nice looking proof right there, thanks!

OpenStudy (anonymous):

could you maybe also help me with this problem too? http://openstudy.com/#/updates/4ed5b89ae4b0bcd98ca8bb86